OpenStudy (anonymous):

integrate dy/(1-y^2)

hartnn (hartnn):

do you know about partial fractions ?

OpenStudy (amistre64):

1+y^2+y^6 ..... ----------------- 1 - y^2 | 1 (1 - y^2) y^2 (y^2 - y^4) y^4 (y^4 - y^6) .... \[\int~\sum y^{2n}~dy=\sum\frac{1}{2n+1}y^{2n+1}\]

OpenStudy (amistre64):

2n is bad ... got confused on the 2468 parts poping out

hartnn (hartnn):

i am so lost lol whay is all that needed ?

OpenStudy (amistre64):

its not needed, but it is a way to approach it :)

OpenStudy (amistre64):

the long division process creates a polynomial representation of it; which can be integrated by the power rule

hartnn (hartnn):

1+y^2+y^6 .... is not y^2n ?

OpenStudy (amistre64):

yeah, i caught that afterwards :) needed more MtDew

hartnn (hartnn):

but its always good to know there ia another way to solve any problem :) thanks!

OpenStudy (amistre64):

i was antisipating the y^6 that was coming up vv 1+y^2+y^4 ..... ----------------- 1 - y^2 | 1 (1 - y^2) y^2 (y^2 - y^4) y^4 (y^4 - y^6)

OpenStudy (amistre64):

its the summation representation of a hyperbolic function, which can also be written in log format